# Application of AI-based Models for Online Fraud Detection and Analysis

### v1

@@ 5.3 Recommendations
Datasets. The reliance on older, established datasets for training AI-based models is a double-edged sword. While they offer a solid foundation for model development and are used for omparative analysis, their static nature may limit their effectiveness in detecting evolving or emerging fraud trends, hence limiting their effectiveness for detecting new types of fraud. Therefore, there is a strong case for incorporating more dynamic and diverse data sources. Recent studies that use custom crawlers to gather data from a variety of online platforms that focus on on a range of fraud types exemplify best practices in this area. These approaches provide real-time, relevant data that can significantly improve the adaptability and accuracy of detection models. Going forward, it is recommended that researchers consider combining established datasets with freshly collected data to create more robust and resilient AI models.
@@
Methodologies Used. Overall the majority of works used stand-alone machine learning and deep learning models for the detection of online fraud. In many cases, NLP use for feature extraction was under-reported or ignored. Work on online fraud activities that involve textual data, should utilize more sophisticated NLP techniques such as transformer-based models (e.g., BERT, GPT) for deeper semantic understanding and better context handling. Although LLMs on their own have their limitations such as generating hallucinated or inconsistent results, they are extremely powerful for context extraction.
@@
Recent works demonstrate hybrid models, where they combine different AI and NLP methodologies to leverage their strengths, which is performing very well. Most existing studies use supervised machine learning models that require labeled data to detect fraudulent activities. Due to challenges in obtaining new labeled data, researchers often rely on existing datasets which may not capture the content of new techniques and tactics employed by scammers. To address these challenges, further exploration of active learning, semi-supervised learning and anomaly-based models that rely on small amount of labelled data or no labelled data is needed. For example, unsupervised or semi-supervised anomaly detection techniques could be studied to identify outliers and novel fraud patterns that may not be present in the training data. Finally, we observed that almost non of these models had real-time applications. There should be a shift in focus where works attempt to optimize models for real-time processing to ensure timely detection and mitigation of fraudulent activities.
@@
Model Performance Reporting. While going through the detected literature for this SLR, we observed many works that were only reporting a subset of performance metrics of their models and frequently relying on accuracy. However, using accuracy on its own especially when the dataset is unbalanced can be misleading. In a dataset where one class has more observations than another (for example having fewer phishing emails compared to not-phishing emails), a model could achieve very high accuracy simply by predicting the majority class (i.e. not-phishing emails) without doing a good job on detecting the phishing emails.
@@
Overall, it is essential that the researchers report a more comprehensive range of performance metrics beyond accuracy alone. This should include precision, recall, F1-score, AUC score or ROC curve. These metrics provide a more complete and nuanced picture of a model’s performance, especially in dealing with imbalanced datasets. In addition, there is a need to conduct and report detailed error analysis to identify common failure cases and the reasons behind them. This can help in understanding the limitations of the model and areas for improvement. Finally, models need to be cross-validated to ensure the robustness of the reported performance metrics. For example, reporting results from multiple folds of data samples can provide a more reliable estimate of the model’s performance.
@@
Reproducibility. Many of the studies failed to clearly explain key and critical aspects of their model development. These include the features engineered and selected, methods used for extracting features, the data used, size of the dataset, partition of the data into training-test sets, and the hyper-parameters used for tuning and training the models.
@@
To this end, we recommend researchers provide access to the code, datasets, and pre-trained models used in their studies through platforms like GitHub, GitLab, or institutional repositories,. This would help improve the reproducibility of their work. Researchers should also ensure that the methodology section is sufficiently detailed to allow others to replicate their study and model. This should include a clear description of pre-processing steps, feature engineering and selection, model hyperparameters, the training-testing data split and training protocols. Furthermore, the use of standardized frameworks and libraries for model implementation (e.g., TensorFlow, PyTorch) could improve reproducibility. Providing comprehensive documentation and setup instructions will further help others understand and reproduce the work easily.
@@
Usability. Most of the studied papers are proof-of-concept, and as a result, the usability of AI-based models has not been addressed. The effective application and use of AI-based approaches depend on successful usability studies that enable users to develop these models into toolkits and provide user feedback. Usability goals are generally determined by efficiency, effectiveness, engagement, error tolerance and ease of use. It is thus imperative to ensure collaboration between the developers of AI based tools and practitioners. However, while the field of technology usability assessment in front line policing is growing 1894801 ; ZAHABI2018161 , there is a lack of usability studies considering the use of AI in preventive policing, including cybercrimes like online fraud.
@@
Limitation/Bias discussion. The majority of studies do not discuss the limitations of their models or data. Researchers should clearly identify and report the limitations of the study, including any assumptions made, potential biases in the data, and limitations of the methodologies used. The overuse of existing labeled datasets could impact the performance of the models, potentially leading to issues such as over-fitting. However, issues like this are not discussed in the majority of works in our SLR. Researchers should be examining the distribution of different classes and any potential sources of bias in the data used for training and testing. Lastly, there is limited discussion on the generalizability of the models across different datasets, contexts, and evolving fraud patterns. Hence, research studies should test their models on a variety of experiments to evaluate generalization performance of the models.
@@ 6 Conclusion
In this systematic literature review, we have examined a wide range of studies focusing on the detection of various fraudulent activities using AI based models that use techniques from Natural Language Processing including machine learning and deep learning. Our goal was to examine the current state-of-the-art AI-based models used for development and training, investigate the sources of data, and assess how these models are evaluated for effectiveness in analysing and detecting fraudulent activities. Due to limited resources, we have restricted the data collection for the SLR to the years 2019-2024. The studies we identified highlight a focus on a wide range of fraudulent activities, with particular attention given to phishing attacks. However, there is growing interest towards the use of more advanced AI like ChatGPT for creating deceptive content as well as a tool that can be used for scam-baiting.
While significant attention have been given to building classification models that could be used to detect fraudulent activities, particularly with hybrid and advanced NLP techniques and deep learning, including LLMs , there remains a considerable room for improvement. The key AI-model development areas that require attention include performance reporting, reproducibility and transparency. Providing detailed performance reporting will help compare and evaluate different models. Improving reproducibility requires ensuring that the studies can be replicated and there is sufficient content for others to achieve this. Increasing transparency means providing clear information on how the AI-based models work and make decisions. This will help fraud practitioners to interpret and understand the models, and mitigate any biases in AI-based models.
Furthermore, most existing models rely heavily on labelled data and supervised machine-learning techniques. Future studies should give some attention to the application of unsupervised and semi-supervised machine learning for detecting fraud. Similarly, the data sources used for training these models are not suitable for capturing dynamic nature of fraud. Future studies should investigate building AI-based models that are able to capture emerging fraudulent patterns.
Addressing these gaps is crucial for creating more robust, reliable, fair, and ethical AI-based systems to detect fraud. By considering the recommended practices, the research community can better understand, detect, and mitigate fraudulent activities and reduce victimisation, ultimately contributing to a safer and more secure digital environment.
@@
Table 3: Data sources and Detection Methods used for Phishing URL detection. A single asterisk (*) indicates that the data is not publicly available. UNK indicates Unclear details. Empty cells indicate missing values. P: Precision, R: Recall, A: Accuracy, F1: F1 Score, AUC: Area under the Curve.
@@ Acknowledgements
The data extracted from the academic papers included in this review will be released with the full accepted paper.
@@
AP and NT drafted the final manuscript. AP conducted the updated academic literature review, designed and conducted the grey literature review, extracted data from the literature, and analysed and interpreted the results of all aspects of the scoping review. SJ and ED provided substantial feedback on the review process and editing of the document. EM contributed to the coding process. AM and SL contributed to the editing process.

### v2

@@ 6.3 Recommendations
Datasets. The reliance on older, established datasets for training AI-based models is a double-edged sword. While they offer a solid foundation for model development and facilitate the comparative analysis of different models, their static nature may limit their effectiveness in detecting evolving or emerging fraud trends, hence limiting their effectiveness in detecting new types of fraud in the real world. Therefore, a strong case exists for incorporating more dynamic and diverse data sources. Recent studies that use custom crawlers to gather data from various online platforms that focus on a range of fraud types exemplify best practices in this area. These approaches provide real-time, relevant data that can significantly improve the adaptability and accuracy of detection models. Going forward, it is recommended that researchers consider combining established datasets with freshly collected data to create more robust and resilient AI models.
@@
Methodologies Used. Overall, most studies reviewed used stand-alone machine learning and deep learning models to detect online fraud. In many cases, NLP techniques used for feature extraction were under-reported or ignored. Working on online fraud activities that involve textual data should utilise more sophisticated NLP techniques such as transformer-based models (e.g., BERT, GPT) for deeper semantic understanding and better context handling. Although LLMs have limitations, such as generating hallucinated or inconsistent results, they are extremely powerful for context extraction.
@@
Recent research demonstrates the utility of hybrid models, which combine different AI and NLP techniques to leverage their strengths. These hybrid models appear to perform very well. Most existing studies have used supervised machine learning models that require labelled data to detect fraudulent activities. Due to the challenges of obtaining new labelled data, researchers often rely on existing datasets that may not capture the content of new techniques and tactics that scammers employ.
To address these challenges, further exploration of active learning, semi-supervised learning and anomaly-based models that rely on small amounts of labelled data or no labelled data is needed. For example, unsupervised or semi-supervised anomaly detection techniques could be studied to identify outliers and novel fraud patterns that may not be present in the training data. Finally, we observed that almost none of the models reported had real-time applications. There should be a shift in focus where researchers attempt to optimize models for real-time processing to ensure the timely detection and mitigation of fraudulent activities.
@@
Model Performance Reporting. While assessing the literature reviewed in this SLR, we observed many studies that only reported a subset of model performance metrics, and authors frequently relied on the Accuracy performance metric alone. This was the case for all of the online fraud types identified and analysed herein. However, using accuracy on its own, especially when the dataset is unbalanced, can be misleading. In a dataset where one class has more observations than another (for example, having fewer phishing emails compared to not-phishing emails), a model could achieve very high accuracy simply by predicting the majority class (i.e. not-phishing emails) without doing a good job of detecting the phishing emails.
@@
Overall, it is essential that researchers report a more comprehensive range of performance metrics beyond accuracy alone. These should include Precision, Recall, the F1-score, AUC score, or ROC curve. These metrics provide a more complete and nuanced picture of a model’s performance, especially when dealing with imbalanced datasets. In addition, there is a need to conduct and report detailed error analysis to identify common failure cases and the reasons behind them. This can help understand the limitations of the model and areas for improvement. Finally, models need to be cross-validated to ensure the robustness of the reported performance metrics. For example, reporting results from multiple folds of data samples can provide a more reliable estimate of model performance.
@@
Reproducibility. Many studies failed to explain key and critical aspects of their model development. This included the features engineered and selected, methods used for extracting features, the data used, the size of the dataset, the partition of the data into training-test sets, and the hyper-parameters used for tuning and training the models.
@@
To this end, we recommend researchers provide access to the code, datasets, and pre-trained models used in their studies through platforms like GitHub, GitLab, or institutional repositories. This would help improve the reproducibility of their work. Researchers should also ensure that the methodology section of their paper is sufficiently detailed to allow others to replicate their model and critically assess it. This should include a clear description of pre-processing steps, feature engineering and selection, model hyperparameters, the training-testing data split and training protocols. Furthermore, the use of standardized frameworks and libraries for model implementation (e.g., TensorFlow, PyTorch) could improve reproducibility. Comprehensive documentation and setup instructions will help others understand and reproduce the work more easily.
@@
Usability. Most of the papers reviewed were proof-of-concept studies, so the usability of AI-based models has not been addressed. The effective application and use of AI-based approaches depend on successful usability studies that enable users to develop these models into toolkits and provide user feedback. Usability goals are generally determined by efficiency, effectiveness, engagement, error tolerance and ease of use. It is thus also imperative to ensure collaboration between the developers of AI-based tools and practitioners. However, while the field of technology usability assessment in front-line policing is growing 1894801 ; ZAHABI2018161 , there is a lack of usability studies considering the use of AI in preventive policing, including its application to cybercrimes like online fraud.
@@
Bias. The majority of studies do not discuss the limitations of their models or data. Researchers should clearly identify and report such limitations, including any assumptions made, potential biases in the data, and methodologies’ limitations. The overuse of existing labelled datasets could impact the performance of the models, potentially leading to issues such as over-fitting. However, issues like this were not discussed in most of the studies reviewed here. Researchers should examine the distribution of different classes and any potential sources of bias in the data used for training and testing. Lastly, there is limited discussion on the generalisability of the models across different datasets, contexts, and evolving fraud patterns. Hence, researchers should conduct sensitivity experiments to evaluate the generalisation performance of their models.
Limitations. As with every study, ours comes with potential limitations. One that stands out lies in the possibility of missed studies. Although we took great care in designing and refining our search string to capture as many relevant publications as possible, the diversity and rapid growth of literature in this field make it likely that some studies were inadvertently omitted. Furthermore, grey literature sources such as pre-prints may lack consistent metadata, which could have further limited our ability to identify all relevant studies.
Finally, the inherent subjectivity in labelling fraud types and selecting the primary focus of each paper poses a challenge. While efforts were made to standardise the categorisation process via discussions between all authors, bias or misclassification could have introduced slight inconsistencies into the analysis. In more detail, the challenges associated with classifying online fraud have been extensively discussed in various reports and academic papers. Rabitti et al. rabitti2024taxonomy and Eling et al. eling2021cyber explain that the field of cyber risks is rapidly expanding, and many taxonomy-based systems have been proposed. At the same time, the partial taxonomies produced by various bodies have resulted in various reports, often not in harmony with one another, which can lead to many frauds being misclassified or the introduction of grey areas and uncertainty. This lack of uniformity in the academic literature renders the identification of online fraud a challenge that is still to be addressed, calling for further research. Similarly, Cohen et al. cohen2019investigation highlight the lack of consistency across said taxonomies and models, emphasizing the need to understand this risk better.
Moving on to literature and taxonomies produced outside academia, a review from the UK’s National Fraud Authority fraudtypologiesuk explains that online frauds are diverse and can be differentiated further, highlighting the evolving nature of online fraud. That review demonstrates how diverse online fraud can be, taking many forms. It compares how different taxonomies and literature reviews use different umbrella terms to classify specific scams and online fraud. Lastly, a white paper by the UK Police Foundation skidmore24 highlights that “Fraud is daunting in terms of its scale and variety”. The report discusses in detail the various methods adopted by fraudsters, the exploited criminal opportunities, and the experiences of the victims. Notably, the author explains that the online fraud landscape is continuously evolving as fraud methodologies and fraudsters adapt to new technological, social, and commercial opportunities before explaining that online fraud is not defined concretely and is often underreported by victims. We acknowledge that while undertaking this systematic literature review, we also faced and confirmed the above challenges regarding the evolving nature and inherent issue of classifying online fraud, which likely affected how we classified the studies selected for qualitative analysis and, hence, the presentation of our analysis and final results.
@@ 7 Conclusion
In this systematic literature review, we have examined a wide range of studies focusing on the detection of various fraudulent activities using AI-based models and Natural Language Processing techniques. Our goal was to examine the current state-of-the-art models and techniques used for development and training, investigate the sources of data used, and assess how these models are evaluated. Due to resource limitations, we restricted the SLR data collection to 2019-2024. The studies we identified covered a wide range of fraudulent activities, but there was a particular focus on phishing attacks. However, there is growing interest in using more advanced, Generative AI content to create deceptive content and tools that can be used for scam-baiting.
Significant attention has been given to building classification models that could be used to detect fraudulent activities. In particular, hybrid models that combine advanced NLP techniques with deep learning, including LLMs, have been developed. However, there remains considerable room for improvement. The key AI-model development areas that require attention include performance reporting, reproducibility, and transparency. Providing detailed performance reporting will help us to compare and evaluate different models. Improving reproducibility is important and requires researchers to provide sufficient details about what they did and how they did it. Increasing transparency means providing clear information on how the AI-based models work and make decisions. This will help fraud practitioners to interpret and understand the models, and mitigate any biases in AI-based models.
Furthermore, most existing models rely heavily on labelled data and supervised machine-learning techniques. Future studies should give some attention to the application of unsupervised and semi-supervised machine learning for detecting fraud. Similarly, the data sources used for training these models are unsuitable for capturing the dynamic nature of fraud. Future studies should, therefore, investigate building AI-based models that can capture emerging fraudulent patterns and their usability in fraud prevention.
Addressing these gaps is crucial for creating more robust, reliable, fair, and ethical AI-based systems to detect fraud. By considering the recommended practices, the research community can help to better understand, prevent, detect, and mitigate fraudulent activities and reduce victimisation, ultimately contributing to a safer and more secure digital environment.
@@
Here we list tables, based on specific scam types, discussed in the body of this manuscript. For the complete data extracted from the academic works included in this review, please refer to the public repository. [^33] https://osf.io/nrx7y/?view_only=ca1050d48c4c4a969817c6d5f677cb87
Table 3: Data sources and Detection Methods used for Phishing URL detection. A single asterisk (*) indicates that the data is not publicly available. UNK indicates Unclear/Unknown/Unspecified details. Empty cells indicate missing values. P: Precision, R: Recall, A: Accuracy, F1: F1 Score, AUC: Area under the Curve.
@@ Acknowledgements
The data extracted from the academic papers included in this review can be found in this public repository: https://osf.io/nrx7y/?view_only=ca1050d48c4c4a969817c6d5f677cb87.
@@
AP and NT drafted the final manuscript. AP conducted the updated academic literature review, designed and conducted the grey literature review, extracted data from the literature, and analysed and interpreted the results of all aspects of the scoping review. NT contributed to the coding process, and supervised all aspects of the study. SJ and ED provided substantial feedback on the review process and editing of the document. EM contributed to the coding process. AM and SL contributed to the editing process.